Bluegill Company sells 14,800 units at $280 per unit. Fixed costs are $207,200, and income from operations is $1,036,000. Det#11)
Answer #1

(a) Variable cost per unit :-

Variable cost = Sales - fixed cost - income from operations

= (14,800 x $280) - 207,200 - 1,036,000

= $2,900,800

Variable cost per unit = 2,900,800/14,800 = $196

(b) Unit contribution margin :-

  = Sales per unit - variable cost per unit

= 280 - 196 = $84 per unit

(c) Contribution margin ratio :-

= Contribution margin/sales

= (4,144,000 - 2,900,800)/4,144,000

= 30%